OpenStudy (anonymous):

Solve 5(^x^2 - ^3x) = 1/25 a. 1,2 b. 1,4 c. 2,2 d. 2,4 e. 1,3 f. 2,5

OpenStudy (jennychan12):

do u mean \[5(\sqrt{x^2}-\sqrt{3x}) = \frac{ 1 }{ 25 }\] ?

OpenStudy (anonymous):

Nope. Everything is correct except the square root before both of them.

OpenStudy (jennychan12):

do you mean \[5^{x-\sqrt{3x}} = 5^{-2}\]

OpenStudy (jennychan12):

5^-2 is the same as 1/25

OpenStudy (anonymous):

maybe u could distribute the 5 to everything in the parentheses and then subtract 1/25 to bring it to the left side so the whole equation could equal zero..then u could use the quadratic equation...I feel like there's an easier way of doing this but that's the only thing I could come up w/.

OpenStudy (anonymous):

Basically you fill in for x with the two choices. (That's what my teacher said) I tried it and got C as my answer. She told me it was incorrect, and to try again.

OpenStudy (anonymous):

For example: 5(^x^2 - ^3x) = 1/25 5(^2^2 - ^3*2) = 1/25 5(^4 - ^6) = 1/24 5^(-2) = 1/24 0.04 = 1/24 1/24 = 0.04 She said it was incorrect, but it's obviously right.

OpenStudy (anonymous):

Ignore the 1/24. I meant 1/25.

OpenStudy (jennychan12):

what the heck is ^x^2? there's no such notation.

OpenStudy (anonymous):

there's something wrong w/ ur teacher b/c none of the answer choices give u 1/25

OpenStudy (anonymous):

Jennychan12: It's the way my teacher taught me to write those things. Solane12: She said to put 1/25 as a percent - which would be 0.04. Divide 1 by 25 and you get 0.04.

OpenStudy (anonymous):

and i'm not sure if u take advanced math but I haven't seen anything that looks like ^x^2 before

OpenStudy (jennychan12):

i agree with @Solane12

OpenStudy (anonymous):

I put that as another way to write it (because that's how I learned it). It's like this:

OpenStudy (anonymous):

Ohhh..makes sense now..i thought u were doing some weird math.. and the answer is A

OpenStudy (anonymous):

Nope, it's just the way my teacher taught it. I'm home-schooled, and she's not the best teacher. Lol But thank you :)

OpenStudy (jennychan12):

-_- that's what i wrote in the beginning...

OpenStudy (anonymous):

C is wrong because one of the choice is 0.2 and 5^.2=1.379 ..

OpenStudy (whpalmer4):

@jennychan12, you really ought to learn how to use the Equation editor button for these things! \[5^{(x^2-3x)}=1/25\]

OpenStudy (anonymous):

Definitely not what you wrote in the beginning Jenny, but thank you Solane!

OpenStudy (anonymous):

and No problem! :)

OpenStudy (whpalmer4):

Hey, you're the one who needs help, right?

OpenStudy (whpalmer4):

Life lesson: make it easy for others to help you.

OpenStudy (anonymous):

Yeah. I understand it. I just didn't see how I got my answer incorrect before. Then Solane12 explained.

OpenStudy (whpalmer4):

Life lesson #2, read more carefully who said what! Sorry jennychan12!

OpenStudy (whpalmer4):

fallingforamber, what I said to jennychan12 was meant for you. My bad!